My question is how to solve this equation:
((x-4)/(sqrt(x-1))+2(x-1)^(1/2)=0

Answer by richwmiller(17219)

sqrt(x-1) and (x-1)^1/2 are the same thing
(x-4)/sqrt(x-1)+sqrt(x-1)2*sqrt(x-1)/sqrt(x-1)=0
((x-4)+2*(x-1))/sqrt(x-1)=0
multiply by sqrt(x-1)
x-4+2x-2=0
3x-6=0
3x=6
x=2
check
-2+2=0
ok